The local botanical society wants to ensure that the gardens in the town park are properly cared for. The group recently spent $100,000 to plant the gardens. The members want to set up a perpetual fund to provide $100,000 for future replantings of the gardens every year. If interest is 5%, how much money would be needed now to forever pay the cost of planting?

A. $162,293

B. $2,100,000

C. $159,000

D. $259,009

Please explain how you got to the solution. Thank you
Answer #1

Cost per year = $ 100000

r = 5%

Present value of money required = 100000 + 100000/(1 + 5%) + 100000/(1 + 5%)2 + ......

This is an infinite geometric progression (GP) with first term = 100000 and common ratio = 1/1.05

Present value of money required = sum of infinite G.P. = 100000/(1 - 1/1.05) = $ 2100000